# On Singular Integrals and Quantitative Rectifiability in Parabolic Space and the Heisenberg Group

### v1

@@ 1.1. Horizontally antisymmetric kernels
With applications in mind, we identify the analogue of the Riesz transform as the distinguished operator with the class of singular integrals we consider. In both the Heisenberg and parabolic settings the Riesz transform has the property that it is odd in the spatial variables with time held fixed [BHHLN3, CLY1, CM, Ho, LM]. Consequently, we make the following definitions:
@@
Admissible kernels in the Heisenberg group. A smooth function $K:\mathbb{H}\backslash\mathbf{0}$ is an admissible kernel (of dimension $3$) if
@@
where $\nabla_{\mathbb{H}}$ is the horizontal gradient [^4] We shall not use the formula for the horizontal gradient explicitly, but rather some few elementary consequences of its boundedness. Section 3 of [CM] is a useful guide of the role of the horizontal gradient in verifying ‘standard properties’ of singular integral kernels.
@@ 1.2. Vertical $\beta$-numbers and uniform rectifiability
$\beta_{\mu,k}(\mathbf{X},r)^{2}=\inf_{L\in\mathcal{VP}_{k}}\frac{1}{\mu(B(\mathbf{X},r))}\int_{B(\mathbf{X},r)}\Bigl(\frac{\operatorname{dist}(\mathbf{Y},L)}{r}\Bigl)^{2}d\mu(\mathbf{Y})$
@@
In the parabolic setting, this definition was introduced by Hofmann-Lewis-Nystrom [HLN]. In the Heisenberg setting it does not appear to be standard, but we adopt it for our purposes here.
@@ 1.3. The precise result in the parabolic case
Fix $k\in\{1,...,n-1\}$. For $m\in\{k,k+1,k+2\}$, consider the collection of measures $\mu=\mathcal{H}^{m}_{|L}\times\nu$ where $L$ is an $m$-plane and $\nu((t-r,t+r))\approx r^{k+2-m}$ for every $t\in\operatorname{supp}(\nu)$ and $r>0$. If $m\in\{k+1,\min(k+2,n+1)\}$ a measure $\mu$ satisfies:
@@
  • the spatial projection of the support of $\mu$ is $L$, so $\beta_{\mu}(B(\mathbf{X},R))\gtrsim 1$ for all $\mathbf{X}\in\operatorname{supp}(\mu)$. As a consequence, (1.2) fails spectacularly.
@@
is the scaled transportation distance from $\mu$ to $\mathcal{M}_{m}$ in the ball $B(x,r)$. We define $\alpha_{\mu,n+1}(B(\mathbf{X},r))=\infty$ for all $\mathbf{X}\in\operatorname{supp}(\mu)$ and $r>0$.
@@
(1.3) $\int_{0}^{R}\int_{B(\mathbf{X},R)}\mathbf{1}_{\{(\mathbf{X},r)\,:\,\alpha_{\mu,k+1}(\mathbf{X},r)<\varepsilon\text{ or }\alpha_{\mu,k+2}(\mathbf{X},r)<\varepsilon\}}d\mu(\mathbf{X})\frac{dr}{r}\leq CR^{k+2}.$
@@
(1.4) $\int_{0}^{R}\int_{B(\mathbf{X},R)}\mathbf{1}_{\{(\mathbf{X},r)\,:\,\alpha_{\mu,n}(\mathbf{X},r)<\varepsilon\}}d\mu(\mathbf{X})\frac{dr}{r}\leq CR^{n+1}.$
@@
As we have seen, an additional condition that has been imposed upon Ahlfors regularity when studying parabolic problems (for instance in [BHMN]) is time symmetric ADR. Here we remark that time symmetric ADR measures satisfy the $\bf HSDC$} condition, and in fact one only requires ‘one-sided’ symmetry either backwards or forwards in time. For $\mathbf{X}=(X,t)$ and a ball $B(\mathbf{X},r)$, we define $B(\mathbf{X},r)_{-}=\{\mathbf{Y}=(Y,s)\in B(\mathbf{X},r)\,:\,s\leq t\}$. An ADR measure $\mu$ is backwards one-sided ADR if $\mu(B(\mathbf{X},r)_{-})\gtrsim r^{k+1}$ for all $\mathbf{X}\in\operatorname{supp}(\mu)$. If $\mu$ is backwards one-sided ADR, then the condition (1.4) is satisfied [^5] The key point here is that if $\mu\in\mathcal{M}_{n}$, then the time measure $\nu$ must have large gaps in its support, and the backward one-sided ADR condition does not allow this., so we get
@@ 1.4. The proof scheme
There is a difficulty in proving Theorems 1.4 and 1.8 by directly adapting the David-Semmes proof scheme from [DS1]. This difficulty is owing to the vampiric measures associated to the problem. A vampiric measure is a $(k+2)$-Ahlfors regular measure such that for every smooth compactly supported function $\psi$ that is odd in space, $\int_{\mathbb{R}^{n+1}}\psi(\mathbf{Y}^{-1}\cdot\mathbf{X})d\mu(\mathbf{Y})=0$ for every $X\in\operatorname{supp}(\mu)$. In the Euclidean version of this problem studied in [DS1], every vampiric measure is (a constant multiple of) the Hausdorff measure of a plane, but in both the parabolic and Heisenberg cases there are vampiric measures that are strict subsets of a vertical planes. This is also mentioned in [Orp]. As a consequence, the proof scheme in [DS1], which hinges on a bilateral approximation argument for planes, has a limitation in how it can extend. Nazarov, Tolsa and the second author developed an alternative approach in [JNT], using ideas developed in [JN1, JN2, JNRT], and here we implement this basic scheme.
The crucial observation is that to run the algorithms developed in [JNT], one can leverage knowing less about the time component of the measure for more knowledge of the spatial component – in particular by making significant use of the $\alpha$-numbers introduced in Section 2.8. After preliminaries in Sections 2 and 3, we carry out a modification of the cylinder blow-up argument from [JNT]. The conclusion of the cylinder blow-up is that if a Carleson measure condition involving the $\alpha$-numbers implies uniform rectifiability (Theorem 4.3). This result holds in both the parabolic and Heisenberg settings. The second part of the paper (Section 5)) then carries out further analysis of vampiric measures in order to verify the hypothesis of Theorem 4.3.
@@ 2.1. Balls, cubes and cylinders
$B_{r}(\mathbf{X})=\{\mathbf{Y}\in\mathbb{R}^{n+1}:\|XY^{-1}\|<r\}.$
@@
Notice that, in the parabolic case, if $\mathbf{X}_{Q}$ is the center of $Q$, then $Q\subset B_{Q}=B(\mathbf{X}_{Q},(n+1)\ell(Q))$, but this is not the case in the Heisenberg group setting (consider the diameter of the cube $Q((N,0,0),1)$ for large $N$). However, one can compare balls centered at the origin in the Heisenberg group to cubes and cylinders. Indeed, we have, for either group structure we have
(2.1) $C(\mathbf{0},r)\subset Q(\mathbf{0},r)\subset B(\mathbf{0},(n+4)r)$
@@ 2.3. Measures
Throughout the paper, by measure we mean locally finite Borel measure.
@@
To avoid repetition in the statements of theorems, we make the following convention: We will make a fixed choice of constant $C_{0}$ and every $(k+2)$-ADR regular measure in the paper has an ADR constant (at most) $C_{0}$. All constants can depend on $C_{0}$ without further mention.
@@
Suppose $\mu_{n}$ is a sequence of $(k+2)$-ADR measures with constant $C_{0}$ that converges weakly to $\mu$. Then $\mu$ is $(k+2)$-ADR with constant $C_{0}$.
@@ 2.4. Dyadic cubes
Since $(\mathbb{R}^{n+1},\|\,\cdot\,\|,\mu)$ is a space of homogeneous type, following [Chr, DS1, DS2] we decompose $\operatorname{supp}(\mu)$ into “dyadic cubes [^6] It is perhaps a little unfortunate that a dyadic cube need not be a cube.”. To attempt to foreshadow, we will do most of our significant analysis on cubes or cylinders as defined above. Dyadic cubes will be primarily used as a bookkeeping mechanism to isolate scales which are particularly essential to the contribution of the Jones square function. As such, we preserve $Q$ to denote a cube and use a different letter to denote dyadic cubes. (For now $E$ but let’s see if we do this or not….)
@@
  • if $E\in\mathbb{D}_{j}$, we have $2^{j}\lesssim\operatorname{diam}(E)\leq 2^{j+2}$ and $\mu(E)\approx 2^{-j(k+2)}$;
  • if $E\in\mathbb{D}_{j}$, there is a point $\mathbf{Z}_{E}=(Z_{E},\tau_{E})\in Q$ (called the “center” of $Q$) such that $\operatorname{dist}\big(\mathbf{Z}_{E},\operatorname{supp}(\mu)\setminus E\big)\gtrsim 2^{j}$.
@@
The properties of the dyadic grid above imply that, for each $E\in\mathbb{E}(\mu)$, there is a constant $c_{*}$ such that (notice this was false with the cylinder)
@@
We now record a standard reformulation of uniform rectifiability in terms of dyadic cubes:
@@
$\sum_{E\in\mathbb{D}(\mu):E\subset E_{0}}\beta_{\mu}(\mathbf{X}_{E},A\ell(E))^{2}\mu(E)\leq C\mu(E_{0})$
@@ 2.6. The Geometric Littlewood-Paley Condition
As in [DS1], we will find the condition of boundedness of a class of Littlewood-Paley operators to be more convenient than CZOs.
@@
$\sum_{E\in\mathbb{D}(\mu):E\subset E_{0}}\Theta_{\mu,\varphi,A}(E)\lesssim_{A,\varphi}\ell(E_{0})^{k+2}.$
We will use Lemma 2.4 to find Carleson families via the following lemma. This is the basic principle that underpins the papers [JNRT, JNT].
@@
Suppose that $\mu$ is good for all $(k+2)$-dimensional Littlewood-Paley kernels, and that $\mathcal{Q}\in\mathbb{D}(\mu)$ is a family of dyadic cubes with the following property: There exists $\Delta>0$, $A>0$ and a finite family $\mathcal{F}\subset\operatorname{Lip}_{0}^{\operatorname{odd}}(\mathbb{R}^{n+1})$ such that
@@
$\Delta\sum_{E\in\mathcal{Q}:E\leq E_{0}}\mu(Q)\leq\sum_{\varphi\in\mathcal{F}:E\leq E_{0}}\Theta_{\mu,\varphi,A}(E)\lesssim\ell(E_{0})^{k+2}$
@@ 2.8. Alpha Numbers
For $m\in\mathbb{N}$, put $\widetilde{\mathcal{M}}_{m}$ to be the collection of $(k+2)$-ADR measures of the form $\mathcal{H}^{m}|_{{}_{L}}\times\nu$ for a linear $m$-dimensional subspace $L\subset\mathbb{R}^{n}$ and a Borel measure $\nu$ on $\mathbb{R}$.
@@
$\mathcal{M}_{m}:=\bigl\{\text{ Borel measures }\omega(E)=\mu\big(E\cdot\mathbf{X}\big)\;:\;\mu\in\widetilde{\mathcal{M}}_{k},;\mathbf{X}\in\mathbb{R}^{n+1}\bigl\}.$
In the parabolic case, measures in $\mathcal{M}_{k}$ are still product measures, but this is no longer the case in the Heisenberg group.
Given a ball $B=B(\mathbf{X},r)\subset\mathbb{R}^{n+1}$, define
(2.6) $\alpha_{m,\mu}(B)=\frac{1}{r^{k+3}}\inf_{\begin{subarray}{c}\omega\in\mathcal{M}_{m}\end{subarray}}\operatorname{dist}_{Q}(\mu,\omega),$
@@
(2.7) $\operatorname{dist}_{B}(\mu,\omega)=\sup\bigg\{\bigg|\int f(d\mu-d\omega)\bigg|:\operatorname{supp}(f)\subset B_{3r}(\mathbf{X}),\,\|f\|_{lip}\leq 1\bigg\}.$
@@
Fix $M>0$, $m$, and $\mu$ a $(k+2)$-ADR measure. Suppose that, for each $\lambda>0$ the collection $E\in\mathbb{D}(\mu)$ such that $\alpha_{\mu,m}(B(\mathbf{Z}_{E},\ell(E)))>\lambda$ is Carleson. Then, for each $\lambda>0$, the collection $E\in\mathbb{D}(\mu)$ such that $\alpha_{\mu,m}(B(\mathbf{Z}_{E},M\ell(E)))>\lambda$ is Carleson.
@@
$\alpha_{\mu_{\ell},m}(B(\mathbf{X},r))\to\alpha_{\mu,m}(B(\mathbf{X},r)).$

### v2

@@ 1.1. Horizontally antisymmetric kernels
With applications in mind, we identify the analogue of the Riesz transform as the distinguished operator with the class of singular integrals we consider. In both the Heisenberg and parabolic settings the Riesz transform has the property that it is odd in the spatial variables with time held fixed [8, 12, 15, 33, 38]. Consequently, we make the following definitions:
@@
[…]
@@
where $\nabla_{\mathbb{H}}$ is the horizontal gradient [^4] We shall not use the formula for the horizontal gradient explicitly, but rather some few elementary consequences of its boundedness. Section 3 of [15] is a useful guide of the role of the horizontal gradient in verifying ‘standard properties’ of singular integral kernels.
@@ 1.2. Vertical $\beta$-numbers and uniform rectifiability
$\beta_{\mu,k}(\mathbf{X},r)^{2}=\inf_{L\in\mathcal{VP}_{k}}\frac{1}{\mu(B_{r}(\mathbf{X}))}\int_{B_{r}(\mathbf{X})}\Bigl(\frac{\operatorname{dist}(\mathbf{Y},L)}{r}\Bigl)^{2}d\mu(\mathbf{Y}).$
@@
In the parabolic setting, this definition was introduced by Hofmann-Lewis-Nystrom [29] in the codimension one setting. In the Heisenberg setting it does not appear to be standard, but we adopt it for our purposes here.
@@ 1.3. The precise result in the parabolic case
Fix $k\in\{1,...,n-1\}$. For $m\in\{k,k+1,k+2\}$, consider the collection of measures $\mu=\mathcal{H}^{m}_{|L}\times\nu$ where $L$ is an $m$-plane and $\nu((t-r,t+r))\approx r^{k+2-m}$ for every $t\in\operatorname{supp}(\nu)$ and $r>0$. If $m\in\{k+1,\min(k+2,n+1)\}$, such a measure $\mu$ satisfies:
@@
  • the spatial projection of the support of $\mu$ is $L$, so $\beta_{\mu}(\mathbf{X},R)\gtrsim 1$ for all $\mathbf{X}\in\operatorname{supp}(\mu)$. As a consequence, (1.2) fails spectacularly.
@@
is the scaled transportation distance from $\mu$ to $\mathcal{M}_{m}$ in the ball $B(x,r)$. We define $\alpha_{\mu,n+1}(B_{r}(\mathbf{X}))=\infty$ for all $\mathbf{X}\in\operatorname{supp}(\mu)$ and $r>0$.
@@
(1.3) $\int_{0}^{R}\int_{B_{R}(\mathbf{X})}\mathbf{1}_{\{(\mathbf{Y},r)\,:\,\alpha_{\mu,k+1}(\mathbf{Y},r)<\varepsilon\text{ or }\alpha_{\mu,k+2}(\mathbf{Y},r)<\varepsilon\}}d\mu(\mathbf{Y})\frac{dr}{r}\leq CR^{k+2}.$
@@
(1.4) $\int_{0}^{R}\int_{B_{R}(\mathbf{X})}\mathbf{1}_{\{(\mathbf{Y},r)\,:\,\alpha_{\mu,n}(\mathbf{Y},r)<\varepsilon\}}d\mu(\mathbf{Y})\frac{dr}{r}\leq CR^{n+1}.$
@@
As we have seen, an additional condition that has been imposed upon Ahlfors regularity when studying parabolic problems (for instance in [10]) is time symmetric ADR. Here we remark that time symmetric ADR measures satisfy the $\bf HSDC$} condition, and in fact one only requires ‘one-sided’ symmetry either backwards or forwards in time. For $\mathbf{X}=(X,t)$ and a ball $B_{r}(\mathbf{X})$, we define $B_{r}(\mathbf{X})_{-}=\{\mathbf{Y}=(Y,s)\in B(\mathbf{X},r)\,:\,s\leq t\}$. An ADR measure $\mu$ is backwards one-sided ADR if $\mu(B_{r}(\mathbf{X})_{-})\gtrsim r^{k+1}$ for all $\mathbf{X}\in\operatorname{supp}(\mu)$. If $\mu$ is backwards one-sided ADR, then the condition (1.4) is satisfied [^5] The key point here is that if $\nu\in\mathcal{M}_{k+1}$ or $\nu\in\mathcal{M}_{k+2}$, then the time measure $\nu$ must have large gaps in its support, and the backward one-sided ADR condition does not allow this., so we obtain the following corollary.
@@ 1.4. The proof scheme
There is a difficulty in proving Theorems 1.4 and 1.8 by directly adapting the David-Semmes proof scheme from [20]. This difficulty is owing to the vampiric measures associated to the problem. A vampiric measure is a $(k+2)$-Ahlfors regular measure such that for every smooth compactly supported function $\psi$ that is odd in space, $\int_{\mathbb{R}^{n+1}}\psi(\mathbf{Y}^{-1}\cdot\mathbf{X})d\mu(\mathbf{Y})=0$ for every $X\in\operatorname{supp}(\mu)$. In the Euclidean version of this problem studied in [20], every vampiric measure is (a constant multiple of) the Hausdorff measure of a plane, but in both the parabolic and Heisenberg cases there are vampiric measures that are strict subsets of a vertical planes. This is also mentioned in [51]. As a consequence, the proof scheme in [20], which hinges on a bilateral approximation argument for planes, has a limitation in how it can extend. Nazarov, Tolsa and the second author developed an alternative approach in [35], using ideas developed in [37, 36, 34], and here we implement this basic scheme.
The crucial observation is that to run the algorithms developed in [35], one can leverage knowing less about the time component of the measure for more knowledge of the spatial component – in particular by making significant use of the $\alpha$-numbers introduced in Section 2.8. After preliminaries in Sections 2 and 3, we carry out a modification of the cylinder blow-up argument from [35]. The conclusion of the cylinder blow-up is that if a Carleson packing condition involving the $\alpha_{\mu,k}$-numbers is satisfied by a $(k+2)$-ADR measure $\mu$, then $L^{2}(\mu)$-boundedness of SIOs with admissible kernels implies uniform rectifiability (Theorem 4.3). This result holds in both the parabolic and Heisenberg settings. The second part of the paper (Section 5)) then carries out further analysis of vampiric measures in order to verify the hypothesis of Theorem 4.3.
@@ 2.1. Balls, cubes and cylinders
$B_{r}(\mathbf{X})=\{\mathbf{Y}\in\mathbb{R}^{n+1}:\|\bf X\cdot\bf Y^{-1}\|<r\}.$
@@
Notice that, in the parabolic case, if $\mathbf{X}_{Q}$ is the center of $Q$, then $Q\subset B_{Q}=B_{(n+1)\ell(Q)}(\mathbf{X}_{Q})$, but this is not the case in the Heisenberg group setting (consider the diameter of the cube $Q_{1}(N,0,0))$ for large $N$). However, one can compare balls centered at the origin in the Heisenberg group to cubes and cylinders. Indeed, we have, for either group structure we have
(2.1) $C_{r}(\mathbf{0})\subset Q_{r}(\mathbf{0})\subset B_{(n+4)r}(\mathbf{0})$
@@ 2.3. Measures
Throughout the paper, by “measure” we mean locally finite Borel measure.
@@
To avoid repetition in the statements of theorems, we make the following convention: we will make a fixed choice of constant $C_{0}$ and every $(k+2)$-ADR regular measure denoted by $\mu$ in the paper has an ADR constant (at most) $C_{0}$. All constants can depend on $C_{0}$ without further mention.
@@
Suppose $\mu_{j}$ is a sequence of $(k+2)$-ADR measures with constant $C_{0}$ that converges weakly to $\mu$. Then $\mu$ is $(k+2)$-ADR with constant $C_{0}$.
@@ 2.4. Dyadic cubes
Since $(\mathbb{R}^{n+1},\|\,\cdot\,\|,\mu)$ is a space of homogeneous type, following [16, 20, 21] we decompose $\operatorname{supp}(\mu)$ into “dyadic cubes [^6] It is perhaps a little unfortunate that a dyadic cube need not be a cube.”. To attempt to foreshadow, we will do most of our significant analysis on cubes or cylinders as defined above. Dyadic cubes will be primarily used as a bookkeeping mechanism to isolate scales which are particularly essential to the contribution of the Jones square function. As such, we preserve $Q$ to denote a cube and use a different letter to denote dyadic cubes.
@@
  • if $E\in\mathbb{D}_{j}$, we have $2^{-j}\lesssim\operatorname{diam}(E)\leq 2^{-j+2}$ and $\mu(E)\approx 2^{-j(k+2)}$;
  • if $E\in\mathbb{D}_{j}$, there is a point $\mathbf{Z}_{E}=(Z_{E},\tau_{E})\in E$ (called the “center” of $E$) such that $\operatorname{dist}\big(\mathbf{Z}_{E},\operatorname{supp}(\mu)\setminus E\big)\gtrsim 2^{j}$.
  • cubes of $\mu$ have “small boundaries”, i.e. if $E\in\mathbb{D}(\mu)$ and $\tau>0$, then $\mu(\{x\in E:\operatorname{dist}(x,\operatorname{supp}(\mu)\setminus E)\leq\tau\ell(E)\})\leq C\tau^{\omega}\mu(E)$ for some uniform $C,\omega>0$.
@@
The properties of the dyadic grid above imply that, for each $E\in\mathbb{D}(\mu)$, there is a constant $c_{*}$ such that (notice this was false with the cylinder)
@@
We now record a standard reformulation of uniform rectifiability in terms of dyadic cubes.
@@
$\sum_{E\in\mathbb{D}(\mu):E\subset E_{0}}\beta_{\mu}(\mathbf{Z}_{E},A\ell(E))^{2}\mu(E)\leq C\mu(E_{0}).$
@@ 2.6. The Geometric Littlewood-Paley Condition
As in [20], we will find the condition of boundedness of a class of Littlewood-Paley operators to be more convenient than CZOs.
@@
(2.6) $\sum_{E\in\mathbb{D}(\mu):E\subset E_{0}}\Theta_{\mu,\varphi,A}(E)\lesssim C(A,\varphi)\ell(E_{0})^{k+2}.$
The constants $C(A,\varphi)$ will be called the “Littlewood-Paley” constants of $\mu$.
We will use Lemma 2.4 to find Carleson families via the following lemma. This is the basic principle that underpins the papers [34, 35].
@@
Suppose that $\mu$ is good for all $(k+2)$-dimensional Littlewood-Paley kernels, and that $\mathcal{Q}\subset\mathbb{D}(\mu)$ is a family of dyadic cubes with the following property: there exists $\Delta>0$, $A>0$ and a finite family $\mathcal{F}\subset\operatorname{Lip}_{0}^{\operatorname{odd}}(\mathbb{R}^{n+1})$ such that
@@
$\Delta\sum_{E\in\mathcal{Q}:E\subset E_{0}}\mu(E)\leq\sum_{\varphi\in\mathcal{F}:E\subset E_{0}}\Theta_{\mu,\varphi,A}(E)\lesssim\ell(E_{0})^{k+2}$
@@ 2.8. Alpha Numbers
For $m\in\mathbb{N}$, put $\widetilde{\mathcal{M}}_{m}$ to be the collection of $(k+2)$-ADR measures of the form $\mathcal{H}^{m}|_{{}_{L}}\times\nu$ for a linear $m$-dimensional subspace $L\subset\mathbb{R}^{n}$ and a Borel measure $\nu$ on $\mathbb{R}$. We note that $m\in\{k,k+1,k+2\}$ in the case that k<n1k<n-1, and $m\in\{n-1,n\}$ in the case $k=n$.
@@
$\mathcal{M}_{m}:=\bigl\{\text{ Borel measures }\omega(E)=\mu\big(E\cdot\mathbf{X}\big)\;:\;\mu\in\widetilde{\mathcal{M}}_{m},;\mathbf{X}\in\mathbb{R}^{n+1}\bigl\}.$
In the parabolic case, measures in $\mathcal{M}_{m}$ are still product measures, but this is no longer the case in the Heisenberg group.
Given $r>0$ and $\mathbf{X}\in\mathbb{R}^{n+1}$, define
(2.7) $\alpha_{m,\mu}(\mathbf{X},r)=\frac{1}{r^{k+3}}\inf_{\begin{subarray}{c}\omega\in\mathcal{M}_{m}\end{subarray}}\operatorname{dist}_{B}(\mu,\omega),$
@@
(2.8) $\operatorname{dist}_{B}(\mu,\omega)=\sup\bigg\{\bigg|\int f(d\mu-d\omega)\bigg|:\operatorname{supp}(f)\subset B_{3r}(\mathbf{X}),\,\|f\|_{\operatorname{Lip}}\leq 1\bigg\}.$
@@
There are constants $C,c>0$ such that the following holds: Fix $M>0$, $m\in\mathbb{N}$, and $\mu$ a $(k+2)$-ADR measure [^7] We are treating the ADR constant as fixed: $C$ and $c$ depend on the ADR constant $C_{0}$ of the measure.. Suppose that, for some $\lambda>0$ the collection $E\in\mathbb{D}(\mu)$ such that $\alpha_{\mu,m}(\mathbf{Z}_{E},\ell(E))<\lambda$ is Carleson. Then the collection $E\in\mathbb{D}(\mu)$ such that $\alpha_{\mu,m}(\mathbf{Z}_{E},M\ell(E))<c\lambda$ is Carleson.
Similarly, suppose that, for some $\lambda>0$ the collection $E\in\mathbb{D}(\mu)$ such that $\alpha_{\mu,m}(\mathbf{Z}_{E},\ell(E))>\lambda$ is Carleson. Then the collection $E\in\mathbb{D}(\mu)$ such that $\alpha_{\mu,m}(\mathbf{Z}_{E},M\ell(E))>C\lambda$ is Carleson.
@@
$\alpha_{\mu_{\ell},m}(\mathbf{X},r)\to\alpha_{\mu,m}(\mathbf{X},r).$